# Network Alignment with Transferable Graph Autoencoders

### v3

@@ 4.4 Alignment and Complexity analysis
Figure 3: Alignment pipeline for T-GAE.
After learning the powerful T-GAE node embeddings, network alignment is performed by solving the linear assignment problem in (2). An illustration of the assignment is presented in Fig. 3. The node features produced by T-GAE are used to calculate a pairwise distance matrix, followed by the greedy Hungarian algo rithm to predict node correspondences.
To analyze the complexity of our approach we study the 3 main parts of T-GAE: a) The design of the input structural features, b) The message-passing GNN that produces node embeddings, and c) the linear assignment algorithm. The computation of our neighborhood-based structural features is expected to take $\mathcal{O}\left(\left|\mathcal{V}\right|\right)$ in real graphs, as proved in Henderson et al. (2011). The computational and memory complexity of the message-passing GNN is $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c\right)$, and $\mathcal{O}\left(\left|\mathcal{V}\right|c\right)$, where $c$ is the width of the GNN. The computational complexity to align the nodes of the graph is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$ since we are using the suboptimal greedy Hungarian. If we want to optimally solve the linear assignment problem we need to use the Hungarian algorithm that has $\mathcal{O}\left(\left|\mathcal{V}\right|^{3}\right)$ complexity. If we want to process large graphs we can use efficient nearest neighbors algorithms with complexity $\mathcal{O}\left(\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ to perform soft linear assignment. Overall the complexity of T-GAE is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$, or $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c+\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ for large graphs.
Table 2: Summary of Dataset statistics
Graph Matching Celegans (Kunegis, 2013) 453 2,025 2,025 Interactome
Arenas (Leskovec and Krevl, 2014) 1,133 5,451 5,451 Email Communication
Douban (Zhang and Tong, 2016) 3,906 7,215 7,215 Social Network
Cora (Sen et al., 2008) 2,708 5,278 5,278 Citation Network
Dblp (Pan et al., 2016) 17,716 52,867 52,867 Citation Network
Coauthor CS (Shchur et al., 2018) 18,333 81,894 81,894 Coauthor Network
Subraph Matching ACM-DBLP (Zhang and Tong, 2019) 9,872 9,916 39,561 44,808 6,352 Citation Network
Douban Online-Offline (Zhang and Tong, 2016) 3,906 1,118 1,632 3,022 1,118 Social Network
@@ 5 Experiments
The experiments included in this section are designed to answer the following research questions:(1)Can T-GAE generate competing graph matching accuracy while being efficient by utilizing transferability of GNN? (2)Is it empirically evident that node features processed by T-GAE is more robust to perturbations than existing GNN based framework? (3)Is T-GAE robust to different graph matching tasks and real-world noise distributions? (4)Is T-GAE generalizable to networks from various domains with different sizes?
To answer these questions, we conduct graph matching experiments by introducing different levels of edge editing under different perturbation models. The results are included in Section 5.3 and Appendix H. We also conducted sub-graph matching experiments to align two different real-world networks that are partially aligned, which are presented in Section 5.4. For each experiment, we include a collection of bench-marked datasets and competing baselines. We also compare the running time of different algorithms, the results are appended in Appendix F.
For graph matching experiments, we consider graphs without node attributes, and design the input to the GNN models, using $7$ structural features proposed in (Berlingerio et al., 2013). The features include the degree of each node, the local and average clustering coefficient, and the number of edges, outgoing edges, and neighbors in each node’s egonet. This input feature is applied for all GNN-based methods. As a result, the performance of NetSimile, vanilla GAE and WAlign provide measures to assess the benefit of using T-GAE for node embedding.
5.1 Datasets and baselines
@@
Table 2 provides a brief overview of the considered networks. Our comparisons are conducted with 3 categories of baseline methods: (a) GNN based methods: WAlign (Gao et al., 2021), GAE and VGAE (Kipf and Welling, 2016); (b) Graph/Node embedding techniques: NetSimile (Berlingerio et al., 2013), Spectral (Umeyama, 1988), DeepWalk (Perozzi et al., 2014), (Grover and Leskovec, 2016b), GraphWave (Donnat et al., 2018) and LINE (Tang et al., 2015). (c) Optimization based graph matching algorithms: S-GWL (Xu et al., 2019a), ConeAlign (Chen et al., 2020) and FINAL (Zhang and Tong, 2016). Note that LINE, VGAE, DeepWalk, and Node2Vec are omitted from some experiments since they show very poor performance. The reason behind that is that they are not permutation equivariant. GraphWave is also excluded from the sub-graph matching experiment, it could not identify correlated nodes in two different graphs. In the case of graphs without attributes FINAL is equivalent to the popular Isorank (Singh et al., 2008) algorithm. FINAL is omitted in sub-graph matching experiments due to weak performance.
5.2 Model Details
As illustrated in Figure 1, the structure of our proposed encoder consists of two MLPs and a series of GNN layers. The node features are processed by a 2-layer MLP and passed to all the GNN layers. We add skip connections between this MLP layer and all the subsequent GNN layers. The outputs of all GNN layers are concatenated and passed to another 2-layer MLP, followed by a linear decoder to generate the reconstructed graph. The model is optimized end to end by equation 8. Note that one can choose different message passing functions as $f$ and $g$ in Equation (4). Our reported results are based on GIN(Xu et al., 2019b).
@@ 5.3 Graph Matching Experiments
To test the performance of the competing methods, we first attempt to match the graphs of Table 2 with permuted and perturbed versions of them. In particular, let $\mathcal{G}$ be a graph of Table 2 with adjacency matrix $\bm{S}$. For each graph we produce 10 permuted-perturbed versions according to $\hat{\bm{S}}=\bm{P}\left(\bm{S}+\bm{M}\right)\bm{P}^{T}$, where $\bm{M}\in\{-1,0,1\}^{N\times N}$ and $\bm{P}$ is a permutation matrix. For each perturbation level $p\in\{0,1\%,5\%\}$, the total number of perturbations is defined as $p|\mathcal{E}|$, where $|\mathcal{E}|$ is the number of edges of the original graph. We consider 2 perturbation models for the probability to edit each edge/non-edge:
  • Uniform perturbation model: In this model we edit the graph by randomly adding and removing edges, the total number of perturbations is defined as $p|\mathcal{E}|$, where $|\mathcal{E}|$ is the number of edges of the original graph, and $p$ is the perturbation level. Each edge and non-edge shares the same probability of being removed or augmented.
  • Degree perturbation model: In this model we only remove edges. Edges with higher degrees are more likely to be removed to preserve the structure of the graph. Specifically, the probability of removing edge $(i,j)$ is set to $\frac{s_{ij}d_{i}d_{j}}{\sum_{ij}{s_{ij}d_{i}d_{j}}}$, where $d_{i}$ is the degree of node $v_{i}$ and $s_{i,j}$ is the $(i,j)$ element of the graph adjacency.
We first test the ability of T-GAE to perform large-scale network alignment and transfer across different datasets. To this end, we train T-GAE according to (7), where $\mathbb{S}$ consist of the small-size networks, i.e., Celegans, Arena, Douban, and Cora. Then we resort to transfer learning and use the T-GAE encoder to produce node embedding on (a) perturbed versions of Celegans, Arena, Douban, and Cora, and (b) larger graphs, i.e., Dblp, and Coauthor CS. Note that neither the larger graphs, nor the perturbed versions of the small graphs were considered during training. This is in contrast with all competing baselines that are retrained on every testing dataset. The average and standard deviation of the matching accuracy for 10 randomly generated perturbation samples are presented in Table 3. Note that we report the matching accuracy for Uniform perturbation model, and results for the Degree model are discussed in Appendix H.
@@
Our first observation is that for zero perturbation most algorithms are able to achieve a high level of matching accuracy. This is expected, since for zero perturbation the network alignment is equivalent to graph isomorphism. Furthermore, there is a clear benefit of processing the NetSimile embeddings with TGAE since it offers up to $43.7\%$ performance increase. When some perturbation is added, the conclusions are straightforward. Our proposed T-GAE markedly outperforms all the competing alternatives and shows the desired robustness to efficiently perform network alignment at $1\%$ and $5\%$ perturbation levels, and its performance is consistent across all datasets and perturbation levels.
Regarding the ability of T-GAE to perform large-scale network alignment the results are definitive. T-GAE enables low-complexity training with small graphs, and execution at larger settings by leveraging transfer learning. In particular, it is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}1\%$. It is also the only method that consistently achieves at least $60\%$ accuracy at $5\%$ perturbation. To the best of our knowledge, our experiments on DBLP (Pan et al., 2016) and Coauthor CS (Shchur et al., 2018) are the first attempts to perform exact alignment on networks at the order of $20$k nodes and $80$k edges.
Comparing T-GAE with vanilla GAE, we observe that GAE is not robust to noise. This highlights the benefit of T-GAE in handling the distribution shift brought by the structural dissimilarity between different graphs. We also compare runtime between T-GAE and all competing methods, and the results are presented in AppendixF.
5.4 Sub-graph Matching Experiments
@@
In this subsection, we test the performance of T-GAE in matching subgraphs of different networks that have aligned nodes (nodes that represent the same entities in different networks). For example, in ACM-DBLP data set, the task is to find and match the papers that appear in both citation networks, whereas in social networks like Douban Online-Offline, we aim to identify the users that take part into both online and offline activities. Note that this is the most realistic graph matching experiment we can perform since we match 2 different real graphs with partially aligned nodes.
This set of experiments is important in proving the robustness and generalizability of the proposed T-GAE framework for the following reason: (1) As we can observe in Table 3, most optimization methods work well on matching the clean graphs. However, their optimization objective usually prevents them from matching graphs with different numbers of nodes, which is the most realistic scenario in real-world applications. (2) Matching two different graphs is considered challenging for the unknown noise distribution, especially for unsupervised approaches.
@@
We observe a significant improvement in matching accuracy with GNN-based methods (TGAE and WAlign) compared to traditional graph or node embedding techniques. These results demonstrate the ability of GNNs to generate expressive and robust node embeddings compared to classical algorithms. In particular, our proposed framework, T-GAE, consistently achieves the best performance among all competing methods. This suggests that the encoder model illustrated in Fig. 1 and the training framework (8) illustrated in Fig. 2, provide an efficient approach to generate powerful node embedding, that is robust to real-world noise distributions for the task of network alignment.
@@
Limitations: Although our approach achieves state-of-the-art performance in aligning real-graphs, approaching network alignment with a learning method, remains a heuristic and does not offer optimality guarantees. Furthermore, in order to process large graphs we cast network alignment as a self-supervised task. As a result in small-scale settings where the task can be tackled with computationally intensive efficient methods, our algorithm is not expected to perform the best. Finally, for large graphs the complexity of T-GAE $\mathcal{O}(|\mathcal{V}|^{2})$ is limiting and therefore our alternative method with complexity $\mathcal{O}(|\mathcal{V}|c^{2}+|\mathcal{E}|c+|\mathcal{V}|\text{log}(|\mathcal{V}|))$ has to be employed.
@@ 6 Conclusion
We proposed T-GAE, a generalized transferable graph autoencoder to perform network alignment. T-GAE can be trained with multiple graphs and produce robust and permutation equivariant embeddings tailored to network alignment. The produced embeddings are related to the spectral decomposition of the graph and are at least as good in graph matching as certain spectral methods. The proposed approach leverages transfer learning and data augmentation and achieves high levels of matching accuracy for graphs with more than $15,000$ nodes. Experiments with real-world benchmarks on both graph matching and subgraph matching tasks demonstrated the effectiveness and limits of the proposed approach.
Appendix A Notation and further reading
@@
Graph matching, as an important problem, has been studied not only by the community of data mining, it has also been mathematically and statistically investigated. A number of approaches have been proposed to solve this problem for Erdős Rényi random graphs $G(n,\frac{d}{n})$. It has been proved that a perfectly true vertex correspondence can be recovered in polynomial time with high probability [Ding et al., 2020]. Furthermore, a sharp threshold has been proved for both Erdős Rényi model and Gaussian model [Wu et al., 2022]. Most recently, a novel approach to calculate similarity scores based on counting weighted trees rooted at each vertex has been proposed Mao et al. [2023]. Such approach has been proved to be effective in solving the aforementioned network alignment problem on random graphs with high probability. Readers are encouraged to refer to the authors of these publications [Ding et al., 2020, Wu et al., 2022, Mao et al., 2023] for further reading.
Appendix B Spectral characterization of GNNs

### v4

@@ 4.3 Alignment and Complexity analysis
Figure 3: The pre-trained encoder operates on out-of-distribution samples. The generated node embeddings are then matched greedily.
After learning the powerful node embeddings, network alignment is performed by solving the linear assignment problem in (2). An illustration of the assignment is presented in Fig. 3. The node features produced by T-GAE are used to calculate a pairwise distance matrix, followed by the greedy Hungarian algo rithm to predict node correspondences. To analyze the complexity of our approach we study the 3 main parts of T-GAE: a) The design of the input structural features, b) The message-passing GNN that produces node embeddings, and c) the linear assignment algorithm.
The computation of our neighborhood-based structural features is expected to take $\mathcal{O}\left(\left|\mathcal{V}\right|\right)$ in real graphs, as proved in Henderson et al. (2011). The computational and memory complexity of the message-passing GNN is $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c\right)$, and $\mathcal{O}\left(\left|\mathcal{V}\right|c\right)$, where $c$ is the width of the GNN. The computational complexity to align the nodes of the graph is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$ since we are using the suboptimal greedy Hungarian. If we want to optimally solve the linear assignment problem we need to use the Hungarian algorithm that has $\mathcal{O}\left(\left|\mathcal{V}\right|^{3}\right)$ complexity. If we want to process large graphs we can use efficient nearest neighbors algorithms with complexity $\mathcal{O}\left(\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ to perform soft linear assignment. However, this efficient algorithm only works to match graphs with its permuted samples. We include detailed discussion in Appendix Section H. Overall the complexity of T-GAE is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$, or $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c+\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ for un-perturbed samples.
@@ 5 Experiments
5.1 Experiments Setup
The experiments included in this section are designed to answer the following research questions:(1) Can T-GAE generate competing graph matching accuracy on real-world networks from various domains with different sizes, while being efficient by utilizing transferability of GNN for large-scale our-of-distribution graphs? We answer this question in Section 5 and Appendix Section G by comparing the performance on matching small to middle sized graphs with unseen perturbed samples, and large scale out-of-distribution networks under different perturbation distributions. (2) Is T-GAE robust to different graph matching tasks and real-world noise distributions? In Section 5.3, we conduct sub-graph matching experiments to align two different real-world networks that are partially aligned. (3) How does the proposed network architecture (Figure 1) and the training objective (Figure 2) contribute to network alignment? To demonstrate the contribution of each proposed component, we conduct ablation studies in Section 5.4, to compare T-GAE with untrained T-GAE, T-GAE trained with GAE objective, and GAE. (4) How much efficiency does T-GAE offer compared to the existing optimization and GNN approaches, and what are the possible trade-offs between efficiency and matching accuracy of T-GAE? We compare the efficiency of different matching algorithms and empirically prove that (a) The matching accuracy of T-GAE can be further improved by leveraging the exact Hungarian algorithm. (b) The efficiency of T-GAE to match graphs with their permuted samples can be enhanced if we adopt the more efficient matching algorithm introduced in Section 4.3. This set of experiments are included in Appendix Section H.
@@ 5.1 Experiments Setup
For each of the above mentioned experiments, we compare T-GAE with three categories of graph matching approaches: (a)GNN based methods: WAlign (Gao et al., 2021), GAE and VGAE (Kipf and Welling, 2016); (b)Graph/Node embedding techniques: NetSimile (Berlingerio et al., 2013), Spectral (Umeyama, 1988), DeepWalk (Perozzi et al., 2014), (Grover and Leskovec, 2016b), GraphWave (Donnat et al., 2018) and LINE (Tang et al., 2015). (c)Optimization based graph matching algorithms: S-GWL (Xu et al., 2019b), ConeAlign (Chen et al., 2020) and FINAL (Zhang and Tong, 2016). Note that LINE, VGAE, DeepWalk, and Node2Vec are omitted from some experiments since they show very poor performance. The reason behind that is that they are not permutation equivariant. GraphWave is also excluded from the sub-graph matching experiment, it could not identify correlated nodes in two different graphs. In the case of graphs without attributes FINAL is equivalent to the popular Isorank (Singh et al., 2008) algorithm, and FINAL is omitted in sub-graph matching experiments due to weak performance.
We include detailed descriptions of our included datasets, implementation details of T-GAE and all the competing baselines in Appendix Section E.
@@ 5.2 Graph Matching Experiments
In this subsection we compare the performance of T-GAE with all competing baselines to match the graphs with permuted and perturbed versions of them. In particular, let $\mathcal{G}$ be a graph with adjacency matrix $\bm{S}$. We then produce 10 permuted-perturbed versions according to $\hat{\bm{S}}=\bm{P}\left(\bm{S}+\bm{M}\right)\bm{P}^{T}$, where $\bm{M}\in\{-1,0,1\}^{N\times N}$ and $\bm{P}$ is a permutation matrix. For each perturbation level $p\in\{0,1\%,5\%\}$, the total number of perturbations is defined as $p|\mathcal{E}|$, where $|\mathcal{E}|$ is the number of edges of the original graph.
Specifically, we train T-GAE according to (7), where $\mathbb{S}$ consist of the small-size networks, i.e., Celegans, Arena, Douban, and Cora. Then we resort to transfer learning and use the T-GAE encoder to produce node embedding for perturbed versions of (a) Celegans, Arena, Douban, and Cora, and (b) larger graphs, i.e., Dblp, and Coauthor CS. Note that none of these perturbed versions were considered during training. This is in contrast with all GNN baselines that are retrained on every pair of networks in the testing dataset. We report the average and standard deviation of the matching accuracy for 10 randomly generated perturbation samples under uniform edge editing in Table 1, where each edge and non-edge shares the same probability of being removed or augmented. We report the results for removing edges according to degrees and the relevant discussed in Appendix G.
@@
T-GAE framework results in a robust and transferable GNN to perform network alignment at a large scale. Our first observation is that for zero perturbation most algorithms are able to achieve a high level of matching accuracy. This is expected, since for zero perturbation the network alignment is equivalent to graph isomorphism. On the smaller networks(Celegans, Arenas, Douban, Cora), T-GAE performs at least as well as the current state-of-the-art optimization approaches (S-GWL and ConeAlign). Specifically, it achieves up to 38.7% and 44.7% accuracy increase compared to S-GWL and ConeAlign, respectively. Regarding the ability of T-GAE to perform large-scale network alignment the results are definitive. T-GAE enables low-complexity training with small graphs, and execution at larger settings by leveraging transfer learning, and it consistently outperforms all competing baselines on the two networks with more than 10k nodes. In particular, it is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}1\%$. It is also the only method that consistently achieves at least $60\%$ accuracy at $5\%$ perturbation. To the best of our knowledge, our experiments on DBLP (Pan et al., 2016) and Coauthor CS (Shchur et al., 2018) are the first attempts to perform exact alignment on networks at the order of $20$k nodes and $80$k edges.
The benefits of processing structural node features with T-GAE is clear. There is a clear benefit of processing the structural embeddings with TGAE since it offers up to $43.7\%$ performance increase compared to NetSimile. When some perturbation is added, the conclusions are straightforward. Especially when perturbations are added, our proposed T-GAE markedly outperforms all the competing GNN alternatives(WAlign and GAE) and shows the desired robustness to efficiently perform network alignment. We observe that neither GAE nor WAlign is robust to noise. This highlights the benefit of T-GAE in handling the distribution shift brought by the structural dissimilarity between different graphs.
5.3 Sub-graph Matching Experiments
@@
We further test the performance of T-GAE in matching subgraphs of different networks that have aligned nodes (nodes that represent the same entities in different networks). Specifically, in ACM-DBLP, the task is to match the papers that appear in both citation networks; in Douban Online-Offline, we aim to identify the users that take part into both online and offline activities. Note that this is the most realistic graph matching experiment we can perform since we match 2 different real graphs with partially aligned nodes.
@@
T-GAE uniformly achieves the best performance in the most realistic scenario of network alignment. Most optimization based approaches do not generalize to this real-world scenario because their optimization objective usually prevents from matching graphs with different numbers of nodes. There is a significant improvement in matching accuracy with GNN-based methods (TGAE and WAlign) compared to traditional graph or node embedding techniques. In particular, T-GAE consistently achieves the best performance among all competing methods. This suggests that the encoder model illustrated in Fig. 1 and the training framework (7) illustrated in Fig. 2, provide an efficient approach to generate powerful node embedding, that is robust to real-world noise distributions for the task of network alignment, compared to the existing GNN frameworks.
5.4 Ablation study
Figure 4: Graph matching performance comparison of T-GAE,T-GAE trained on a single graph(T-GAE single), the untrained T-GAE (T-GAE untrained), and GAE. The proposed training objective and encoder structure helps to prompt the expressiveness of GNN thus achieve higher accuracy as we introduce more perturbations.
The proposed architecture and training objective prompts the robustness of GNN when matching graphs with their highly perturbed versions. From Figure 4, we observe that T-GAE outperforms the untrained T-GAE by a great margin when matching highly permuted samples. This implies that the proposed training objective effectively improves the robustness of GNN, which is the key property to deploy GNN to match perturbed graphs. Further, the performance gap between GAE and T-GAE single underscores the efficacy of incorporating attention mechanism on each layer for both input and output node features, as illustrated in Figure 1.
6 Limitation
@@
Although our approach achieves state-of-the-art performance in aligning real-graphs, on both graph matching and sub-graph matching tasks, approaching network alignment with a learning method, remains a heuristic and does not offer optimality guarantees. Furthermore, in order to process large graphs we cast network alignment as a self-supervised task. As a result in small-scale settings where the task can be tackled with computationally intensive efficient methods, our algorithm is not expected to perform the best. Finally, the complexity of T-GAE $\mathcal{O}(|\mathcal{V}|^{2})$ is limiting, this bottleneck comes from the greedy linear assignment algorithm to match the node embedding, and therefore the alternative method with complexity $\mathcal{O}(|\mathcal{V}|c^{2}+|\mathcal{E}|c+|\mathcal{V}|\text{log}(|\mathcal{V}|))$ should be deployed when we match very large scale graphs with their permuted versions.
@@ 7 Conclusion
We proposed T-GAE, a graph autoencoder framework that utilizes transferability and robustness of GNN to perform network alignment. T-GAE is an unsupervised approach that tackles the high computational cost of existing optimization based algorithms, and can be trained on multiple small to middle sized graphs to produce robust and permutation equivariant embeddings for larger scale unseen networks. We proved that the produced embeddings of GNNs are related to the spectral decomposition of the graph and are at least as good in graph matching as certain spectral methods. Our experiments with real-world benchmarks on both graph matching and sub-graph matching demonstrated the great potential of utilizing the good properties of GNNs to solve network optimization problems in a more efficient and scalable way.
Appendix A Appendix
Appendix B Notation
@@
Appendix C Network Alignment in broader domains
In this paper, we study the general form of network alignment to match nodes between two graphs. This task is challenging because of the little given information. However, it’s not trivial to correctly and efficiently utilize the information contained in text-rich networks, such as KGs [He et al., 2024]. In the domain of knowledge graph (KG) mining, entity alignment (EA) is the task to identify the same entities existing in knowledge graphs. [Ge et al., 2021a] proposes to replace the labor-intensive pre-processing with entity names mining. A structural-based refinement procedure is then applied to refine the entity name matching results. [Ge et al., 2021b, Gao et al., 2022] solves large scale EA by aligning the KG in mini-batches. [Ge et al., 2023] further proposes a self-supervised EA framework by automatically generating positive and negative matched node pairs. [Liu et al., 2023b] generalizes the unsupervised matching algorithm to temporal KGs. Specifically, it encodes the temporal and relational information respectively before an innovative jointly decoding process. Recently, the ability to deploy EA algorithms in real-world scenarios is enhanced by [Zhu et al., 2023], which aligns multi-modal knowledge graphs.
Network alignment, as an important problem, has been studied not only by the community of data mining, it has also been mathematically and statistically investigated. A number of approaches have been proposed to solve this problem for Erdős Rényi random graphs $G(n,\frac{d}{n})$. It has been proved that a perfectly true vertex correspondence can be recovered in polynomial time with high probability [Ding et al., 2020]. Furthermore, a sharp threshold has been proved for both Erdős Rényi model and Gaussian model [Wu et al., 2022]. Most recently, a novel approach to calculate similarity scores based on counting weighted trees rooted at each vertex has been proposed Mao et al. [2023]. Such approach has been proved to be effective in solving the aforementioned network alignment problem on random graphs with high probability. Readers are encouraged to refer to the authors of these publications [Ding et al., 2020, Wu et al., 2022, Mao et al., 2023] for further reading.
Appendix D Proof of Theorem 3.2
D.1 Spectral characterization of GNNs